Find the product of 2x.5x.6x using the law of exponent?
1
Expert's answer
2015-01-12T04:28:34-0500
The problem can be rewritten in such way:

2*5*6*x*x*x

1st step
2*5*6=60

2nd step
When we multiply the three x's together, each x is being raised to the 1st power so it can be rewritten

(x^1)*(x^1)*(x^1)

According to the laws of exponents, when we multiply numbers that have the same bases, we can simply add the exponents, so we add the three 1's together to get three:

(x^1)*(x^1)*(x^1)=x^3

ANSWER:60x^3